Use the production function:Y=K^(3/4)*L^(3/4) to Show that marginal product is decreasing for both capital and labor(5pts), yet there are increasing returns to scale(5pts).
Answer #1

MPl=dY/dL = 3/4 K3/4 L3/4-1

MPl =3/4 K3/4 / L1/4

MPl and L are inversely related.

MPk=dY/dK = 3/4 L3/4 K3/4-1

MPk =3/4 L3/4 / K1/4

MPk and K are inversely related.

It means with an increase in input their marginal product decreases.

It is a Cobb Douglas production function. So, we can find out return of scale by adding powers of input in the production function.

SUm of powers = 3/4 + 3/4 = 1.5 Sum is more than 1 and hence it is increasing returns to scale means increase in the input will increase output at an increasing rate.